### v1

# The Energy Cost of Reasoning: Analyzing Energy Usage in LLMs with Test-time Compute

Abstract
Scaling large language models (LLMs) has driven significant advancements, yet it faces diminishing returns and escalating energy demands. This work introduces test-time compute (TTC)—allocating additional computational resources during inference—as a compelling complement to conventional scaling strategies. Specifically, we investigate whether employing TTC can achieve superior accuracy-energy trade-offs compared to simply increasing model size. Our empirical analysis reveals that TTC surpasses traditional model scaling in accuracy/energy efficiency, with notable gains in tasks demanding complex reasoning rather than mere factual recall. Further, we identify a critical interaction between TTC performance and output sequence length, demonstrating that strategically adjusting compute resources at inference time according to query complexity can substantially enhance efficiency. Our findings advocate for TTC as a promising direction, enabling more sustainable, accurate, and adaptable deployment of future language models without incurring additional pretraining costs.
1 Introduction
Test-time compute (TTC) has emerged as a promising approach to significantly enhance the performance of large language models (LLMs) without the reliance on traditional scaling laws. Conventionally, improving neural language models involved scaling either the number of parameters or expanding the training dataset [2020arXiv200108361K]. However, this approach demands substantial computational resources, resulting in considerable energy consumption during both training and inference phases. Furthermore, such scaling is increasingly constrained by the limited availability of sufficiently large and diverse datasets[2022arXiv221104325V], highlighting diminishing returns in terms of resource efficiency.
Recent research addresses this limitation by proposing a new strategy to enhance the performance of LLMs: allocating additional computational resources at inference time, known as TTC. TTC is inspired by human reasoning process, where humans reflect and reason logically to answer challenging questions rather than responding impulsively. Similarly, LLMs should leverage additional computational effort during inference to improve their responses. Prominent examples of TTC methodologies include prompting techniques like Chain-of-Thought (CoT) reasoning  [2022arXiv220311171W, 2023arXiv230510601Y], self-revision methods [2024arXiv240601297K, 2024arXiv240615673L, 2023arXiv230317651M], post-training for reasoning [deepseek, star, rest], and ensemble-like approaches aggregating response from multiple parallel generations to improve the final response [llmonk, getting_50, 2023arXiv230306135I, 2022arXiv220311171W]. Despite methodological differences, these strategies collectively emphasize additional computational investment during inference to enhance model accuracy.
Nevertheless, the rising computational demands of TTC necessitate careful consideration of associated energy costs. Recent industry reports underscore this concern, with Meta attributing up to 70% of its AI power consumption to inference processes [wu2022sustainable], Google reporting 60% of machine learning energy usage [patterson2022carbon], and AWS indicating inference-related demands accounting for 80-90% of computing resources [barr2019amazon] all without TTC. Moreover, broader data center energy consumption trends indicate a rapid escalation, expected to constitute up to 12% of total U.S. electricity usage by 2028 [shehabi2024united]. The environmental impact and financial implications of such intensive energy consumption underline the urgency of evaluating the sustainability and efficiency of TTC approaches.
In this paper, we critically examine the performance-energy trade-offs associated with TTC compared to traditional model scaling. Specifically, we define performance as accuracy of evaluating benchmarks and investigate the relationship between accuracy improvements and energy consumption incurred by employing TTC methods versus scaling model sizes. Through comprehensive experiments conducted on Qwen2.5 [qwen2.5] models (1.5B, 7B, 14B, and 32B parameters) using A100 GPUs across mathematical, coding, and common-sense reasoning benchmarks, we derive key insights into the comparative advantages and costs of TTC.
Our findings highlight several crucial observations:
  • TTC demonstrates superior accuracy-energy trade-off improvements over traditional model scaling especially for performant models. Notably, smaller models enhanced with TTC can outperform substantially larger models relying solely on scale.
  • However, indiscriminate application of TTC can drastically elevate energy consumption, with cases where TTC usage amplifies energy costs up to 97.48$\times$, as observed with the 7B model on a coding benchmark.
  • Output sequence length emerges as a reliable indicator of model comprehension, with extended sequences often signaling models’ struggles or attempts at multiple reasoning pathways, consequently incurring higher energy usage.
  • Complex questions inherently drive models to allocate greater computational resources during inference, paralleling human cognitive processes in handling challenging tasks.
To the best of our knowledge, this study represents one of the first systematic explorations into the energy implications of TTC for LLM inference. Our analysis provides a foundational understanding of how TTC can serve as a more energy-efficient alternative and complement to traditional scaling while also underscoring the potential for substantial resource expenditure. We explicitly focus on inference-related impacts, leaving exploration of training-phase considerations as important avenues for future research.
Figure 1: Accuracy vs Energy per query. Each dot in the line represent Qwen2.5 1.5B, 7B, 14B, and 32B from left to right, respectively.
2 Background
LLM Inference Pipeline
Inference for transformer-based LLMs [llama, qwen2.5, gemini, gpt4] typically involves two sequential stages: prefill and decode. Initially, when an input sequence is provided, it undergoes the prefill stage, where all tokens in the seqeunce are processed simultaneously. This stage computes the first output token and generates a Key-Value cache (KV-cache), which stores contextual information from the input sequence. Following the prefill stage, the model enters the decode stage, wherein it autoregressively generates subsequent tokens one-by-one, leveraging the previously generated KV-cache.
Each of these stages presents distinct computational bottlenecks. The prefill stage is predominantly compute-bound due to extensive matrix multiplications resulting from processing all tokens simultaneously. In contrast, the decode stage typically becomes memory bandwidth (BW)-bound because of the frequent accesses to the KV-cache. Recognizing this distinction, state-of-the-art LLM inference engines optimize latency and throughput by overlapping these two stages [orca, vllm, sglang, nanoflow, deepspeed], efficiently exploiting their differing computational characteristics within the same model architecture.
Test-time Compute
Recently, leveraging additional computational resources at inference time—known as TTC—has emerged as an effective strategy for enhancing the reasoning capabilities and overall accuracy of LLMs. We categorize TTC techniques into two categories based on their resource utilization patterns: 1) techniques that increase the input tokens more [llmonk, 2022arXiv220311171W, few_shot] or 2) those that increase output tokens more [deepseek, 2023arXiv230317651M, star, rest]. Each category uniquely affects the resource demands of the inference pipeline.
Increasing the number of input tokens primarily impacts the prefill stage, significantly intensifying its computational load, as the compute requirement scales quadratically with the length of the input sequence. On the other hand, increasing the number of output tokens predominantly affects the decode stage. Specifically, extending the output length linearly increases both the size of the KV-cache and the number of autoregressive decoding steps, thus elevating memory BW requirements.
Energy Usage Measurement
While considerable research has focused on improving the performance and reasoning capabilities of LLMs, the energy consumption associated with their deployment and usage remains relatively underexplored. Typically, energy consumption is measured by tracking power usage throughout model runtime and integrating this measurement over time.
Early pioneering efforts, such as the study by Strubell et al. [strubell2020energy], began to systematically investigate the energy cost of training transformer models, spurring subsequent studies that quantified energy consumption during model training [patterson2022carbon]. With the increasing deployment of LLMs in practical applications, recent studies have expanded this scope to include energy measurements specific to inference workloads [luccioni2024power, fernandez2025energy, wu2025unveiling, patel2024characterizing, stojkovic2024towards].
On top of the existing works, we specifically examine the additional energy overhead incurred by TTC methods. Despite TTC becoming a widely accepted practice to improve the reasoning capabilities of LLMs, the precise energy costs associated with such approaches remain largely unexplored—highlighting the critical importance and novelty of our analysis.
3 Methodology
Table 1: List of benchmarks used in this study
Task Benchmark Task Benchmark Task Benchmark
Math AIME24 [aime24] Code HumanEval [humaneval] Common Sense HellaSwag [hellaswag]
AIME25 [aime25] LiveCodeBench [livecodebench] MMLU [mmlu]
GPQADiamond [gpqa] MBPP [mbpp] GSM8K [gsm8k]
Math500 [math500] CodeForces [codeforces] CommonsenseQA [talmor-etal-2019-commonsenseqa]
In this study, we systematically measure the additional energy consumption associated with TTC methods in the context of LLM inference. To quantify this energy overhead, we base our work on Evalchemy [evalchemy] and utilize SGLang [sglang] for model execution on NVIDIA A100 GPUs [a100], each with a maximum power rating of 500W. We run each benchmark from start to end assuming that all questions are present to be processed to eliminate any scheduling overhead or queuing delay. Power consumption during model inference is monitored using the NVIDIA Management Library (NVML) [nvml]. We integrate power usage over the inference period to compute total energy consumption for each model evaluation.
Categorization of TTC Methods
We group TTC approaches into two distinct categories: (1) methods that increase the number of input tokens and (2) methods that expand the number of output tokens. For the first category, we select parallel sampling [llmonk, getting_50, 2023arXiv230306135I, 2022arXiv220311171W] with majority vote (MV) to finalize the answer. Although parallel sampling can leverage prefix caching by performing the prefill stage only once, we deliberately opt to repeat the prefill stage for each sample independently. This strategy ensures effective utilization of varying sampling temperatures, thereby producing a diverse set of candidate outputs.
For the second category, we adopt the reasoning token (RT) approach [deepseek, star, rest]. Unlike baseline models that lack post-training enhancements and thus cannot effectively refine their outputs [2023arXiv231001798H], models utilizing reasoning tokens have demonstrated a notable capacity for self-correction and improved reasoning. This enhanced capability results from specialized tokens learned during the post-training phase.
Model Selection
We base our study on models supported by the DeepSeek-R1 framework [deepseek], selecting four representative models from the Qwen2.5 family [qwen2.5]: 1.5B, 7B, 14B, and 32B variants. These models have been distilled from DeepSeek-R1 and are particularly suitable for demonstrating the effectiveness of reasoning tokens in refining model outputs. To facilitate a rigorous comparison, we also employ these same Qwen2.5 models as baseline (Base) references (without TTC) and in parallel sampling scenarios. Using identical model architectures across conditions ensures a fair and controlled evaluation. 32B models are executed across two NVIDIA A100 GPUs in tensor parallel manner as they on average use over 100GB compared to 80GB memory on our A100 GPUs.
Evaluation Tasks
To comprehensively capture the impact of TTC, we evaluate model performance across three representative machine learning task categories, each consisting of four distinct benchmarks. Specifically, we choose mathematical reasoning and code-generation tasks to assess how effectively TTC improves a model’s logical reasoning capabilities. Additionally, common sense tasks are selected to investigate whether extra computation at inference time can meaningfully enhance the model’s retention or representation of factual information. Detailed descriptions of tasks and corresponding datasets are summarized in Table 1.
4 Results
Figure 2: Normalized energy. The first, second, and third four benchmarks represent math, code, and common-sense, respectively. The dotted horizontal line represents the baseline.
We present our results on Base, MV, and RT. Evaluations are run 10 times with batch size 16, maximum sequence length 32768, and data type Bfloat16 unless stated otherwise. We generate 5 samples when running MV with parallel sampling. Common-sense benchmarks take roughly 10 minutes per evaluation while math and code benchmarks take between 1 to 4 hours depending on the benchmark.
4.1 Overview
Accuracy vs Energy per Query
The primary objective of this analysis is to explore the trade-off between accuracy improvements and associated energy costs when employing TTC. Figure 1 presents this trade-off across mathematical, coding, and common-sense benchmarks.
Firstly, the slope of the accuracy-energy curve is steeper for math and common-sense tasks, with a clearly defined knee point situated higher and further from the origin for RT. This indicates a superior accuracy-to-energy ratio when using RT. For example, in mathematical benchmarks, increasing the model size from 1.5B to 7B parameters in the Base configuration yields only a 4.8% accuracy improvement with negligible energy difference per query. In contrast, using RT on the 1.5B model significantly boosts accuracy by 34.3%, representing a 7.5-fold increase in accuracy improvement over traditional scaling, at similar energy consumption levels.
In coding benchmarks, scaling model size offers better accuracy/energy gains at lower model scales than RT. Specifically, transitioning from a Base 1.5B to Base 7B model increases accuracy by 16.8% while reducing energy consumption per query by 40.1%. Conversely, applying RT to the 1.5B model achieves a similar accuracy improvement (17.3%) but increases energy per query by 57.4%. However, when targeting higher accuracy regimes (e.g., Base 32B model versus RT 7B model), RT demonstrates a superior slope of accuracy improvement per unit energy consumed (28.1%/Wh for RT versus 20.7%/Wh for Base). To ensure a fair comparison, RT 7B was chosen to match energy usage with the larger Base model, confirming that RT is more advantageous at higher accuracy demands.
Common-sense benchmarks show limited or even adverse performance when employing RT. Specifically, the Base 1.5B model outperforms both the 1.5B and 7B RT configurations. This is expected because common-sense tasks predominantly evaluate factual knowledge retrieval rather than complex reasoning capabilities, limiting the effectiveness of RT.
Secondly, the use of MV combined with parallel sampling substantially increases energy consumption without proportionate accuracy improvements. This inefficiency occurs because MV relies on simplistic aggregation, where the probability of correctness, $p$, remains unchanged. Particularly, if $p<0.5$ and we sample 5 generations, using MV with multiple candidates can potentially reduce accuracy due to the binomial distribution of generations. Consequently, a more sophisticated aggregation method is necessary to effectively utilize candidate outputs and justify increased energy consumption.
Finally, an intriguing anomaly emerges within the coding benchmarks: Base and MV configurations for the 14B models exhibit lower accuracy compared to their 7B counterparts. This finding reinforces the notion that merely increasing model size does not guarantee enhanced accuracy if underlying reasoning capabilities remain insufficient.
Figure 3: Power readings during runtime.
Normalized Energy
We analyze the energy consumption implications of employing TTC, specifically quantifying the energy increase introduced by MV and RT relative to Base. Figure 2 illustrates these normalized energy metrics.
On average, MV consumes 2.63$\times$, 2.01$\times$, and 4.69$\times$ more energy than the Base configuration for mathematical, coding, and common-sense benchmarks, respectively. RT demonstrates a more substantial average energy increase, using 3.66$\times$, 10.4$\times$, and 1.08$\times$ more energy than Base for math, code, and common-sense tasks, respectively.
The energy cost of RT, in particular, can escalate dramatically, with a peak increase of 97.48$\times$ for the 7B model on the HumanEval coding benchmark. This extreme rise is largely driven by a substantial growth in the number of generated tokens; RT models produce, on average, 4.4$\times$ more tokens than Base, and in extreme cases up to 46.26$\times$ more tokens.
In contrast, although MV multiplies both input and output tokens by a factor of five, the short output sequences do not fully utilize GPU compute and memory BW during decoding. Consequently, this allows MV to maintain minimal latency overhead despite an increased batch size, resulting in less than a proportional 5$\times$ increase in energy consumption. However, common-sense benchmarks, which generate only a single output token correlating the answer to multiple-choice options, consistently exhibit approximately 5$\times$ energy usage relative to Base.
Power Measurements
We explore power consumption implications of employing TTC. Data centers typically have strict power provisioning limits, requiring accurate assessments of power usage to prevent exceeding provisioned capacity. Figure 3 provides average GPU power measurements and associated standard deviations across each benchmark.
Our analysis reveals that TTC only marginally affects power usage. Specifically, MV introduces an average increase of approximately 5% in power consumption due to increased batch sizes from parallel sampling. Conversely, RT tends to use slightly less power than Base. This unexpected reduction occurs because RT inference is predominantly decode- and especially attention layer-bound, characterized by extremely long output sequences. The resulting high memory requirements for the KV cache create a bottleneck that saturates GPU memory BW but significantly under-utilizes the GPU compute units.
For instance, RT benchmarks produce an average of 7845 tokens per query, translating to an average KV cache size of about 258MB, even for relatively small (7B) models. Given the memory BW saturation threshold (approximately 100MB data array for A100 GPUs [9307857]), the inference process becomes memory-bound, limiting the effective utilization of GPU computational resources. Consequently, the power usage stabilizes at a level consistent with these memory constraints.
4.2 In-depth Analysis
We conclude that parallel sampling with MV shows lower accuracy/energy than RT and that TTC has limited benefit on common-sense benchmarks. To this end, we provide a thorough analysis on RT using two reasoning benchmarks one from math (MATH500) and coding (LiveCodeBench) task each.
Figure 4: Energy vs Accuracy per length. Each line is represent a model and dot represents maximum sequence length.
Length Sweep
We investigate how the maximum output sequence length affects RT performance and associated computational costs. By incrementally varying the maximum sequence length by steps of 3277 tokens—one-tenth of the 32768 tokens—we assess the impact on accuracy-energy trade-off.
Figure 4 demonstrates the relationship between accuracy and energy consumption for two representative benchmarks. Notably, RT consistently forms a distinct and superior Pareto frontier compared to Base models. This distinction arises because Base models reach an accuracy ceiling even when allowed extended sequence lengths, with improvements requiring substantially larger model sizes. Base accuracy saturates at 48.3% for MATH500 and 23.5% for LiveCodeBench, whereas RT achieves significantly higher accuracy ceilings of 89.8% and 72.1%, respectively. Moreover, the RT Pareto frontier highlights that slight accuracy trade-offs can yield considerable energy savings, valuable in resource-constrained environments. For example, transitioning from RT 32B to RT 14B to solve LiveCodeBench reduces energy consumption by 1.74× while sacrificing only 7.8% accuracy.
Additionally, RT models often terminate computations before reaching the maximum sequence length, which can be seen from the clustered points at the end of all RT lines. This indicates their ability to avoid excessive processing when uncertain, thus efficiently managing resources.
Figure 3(a) reveals the competitive potential of smaller RT models. For instance, the 1.5B RT model achieves 83.2% accuracy with just 135.22Wh, outperforming the Base 32B model at only 45.7% accuracy and consuming 424.54Wh. Remarkably, the RT 7B model reaches an accuracy of 89.6%, approaching the performance of much larger models such as DeepSeek-R1 with 617B parameters.
Tokens per Correct and Incorrect Queries
Figure 5 illustrates the number of output tokens generated for correct and incorrect answers in both Base and RT models highlighting how token length relates to model reasoning.
Base models exhibit tightly clustered sequence lengths around 200 tokens for both correct and incorrect answers, with numerous outliers among incorrect responses. These models’ limited reasoning capabilities result in minimal token generation for answering, and outlier sequences typically represent unnecessary token generation rather than deeper reasoning. Increasing the model size reduces the number of outliers but maintains clustering near short sequence lengths, underscoring their restricted reasoning abilities.
Conversely, RT models utilize fewer tokens to reach correct answers compared to incorrect responses. For correct queries, token length distributions are consistent across model sizes, suggesting that RT effectively guides reasoning up to a certain point. However, when uncertain, RT models produce longer incorrect responses due to loops or divergent reasoning. Larger models significantly reduce the token lengths of incorrect answers, indicating a stronger capability to recognize ineffective reasoning paths and terminate more efficiently.
Figure 5: Output token count distribution of correct and incorrect queries.
Query Difficulty
We analyze the impact of task difficulty on energy consumption and accuracy using RT. Figure 6 depicts the accuracy versus energy-per-query tradeoff across easy, medium, and hard difficulty levels for selected benchmarks.
Firstly, task difficulty directly correlates with increased token generation and higher energy consumption, with each difficulty level forming distinct Pareto frontiers. This indicates alignment between model-perceived and human-perceived task difficulty, where additional computational effort reflects increased complexity. This observation expands on previous findings that models’ and humans’ perceptions of difficulty align, especially for reasoning-based tasks.
Secondly, for easier tasks, smaller models with RT can match the performance of larger models, resulting in significant energy savings. Specifically, deploying 14B and 1.5B models instead of a 32B model saves 11.55$\times$ and 1.48$\times$ energy for easy tasks in MATH500 and LiveCodeBench, respectively.
Further analysis reveals that using RT provides better accuracy-per-energy improvements than simply increasing model size, even when all models employ RT. For example, on the hard queries of MATH500, the 14B model increases accuracy from 46.2% to 71.6% at an additional 0.4Wh, whereas scaling to a 32B model has marginal impact on accuracy while using 0.6Wh more energy. Similarly, the 14B model for medium-difficulty tasks in LiveCodeBench improves accuracy from 10.0% to 36.2% with an additional 0.96Wh of energy, demonstrating an accuracy/energy benefit of 39.5%/Wh compared to only 3.33%/Wh when scaling up to the 32B model. Thus, RT consistently offers superior or equivalent accuracy-per-energy tradeoffs compared to increasing model sizes.
Figure 6: Accuracy vs Energy per difficulty level.
Impact of Batch Size on Throughput
We examines the influence of batch size on throughput (queries-per-second, QPS) of RT compared to Base. Increasing batch size is a common strategy to enhance throughput, particularly important given the increased latency and reduced throughput associated with the longer output sequences generated by RT.
Figure 7 depicts the QPS across various batch sizes for the two benchmarks using the highest-performing 32B models. Results indicate that larger batch sizes consistently improve throughput for both configurations. For Base, output sequences are relatively short, resulting in smaller KV cache footprints that under-utilize GPU memory BW. Although RT applied to MATH500 generates more tokens than Base, it still generates short enough sequences to under-utilize GPU memory BW. As a result, even at high batch sizes, the KV cache remains manageable, and throughput scales proportionally with batch size increases.
However, RT applied to the LiveCodeBench exhibit a notable saturation in throughput at a batch size of approximately 128. On average, RT generates around 6588 output tokens for LiveCodeBench, shifting inference into a predominantly decode-bound stage. During the decode stage, loading the large KV cache that scales linearly with batch size from GPU memory becomes the critical performance bottleneck, eventually saturating available GPU memory BW.
Figure 7: Throughput vs Batch size.
5 Conclusion and Discussion
In this work, we systematically evaluate the energy consumption of large language model (LLM) inference specifically through Test-Time Computation (TTC). We investigated two prevalent TTC strategies: (1) generating multiple candidate answers, and (2) self-refinement through iterative reasoning. Our analysis reveals that employing TTC strategies generally achieves better accuracy-energy trade-offs than solely increasing model size. Additionally, output sequence length emerges as a reliable indicator of model comprehension and complex questions require greater computational resources during inference. Combining larger models with TTC strategies yields the highest accuracy but comes with substantial energy costs. Thus, a prudent approach is essential, balancing accuracy gains against energy expenditures. Our findings suggest promising directions for leveraging TTC in a more energy-efficient manner.
Difficulty-aware Model Selection
Our results (Figure 4) demonstrate that smaller models with TTC can compete effectively against larger models, with distinct Pareto frontiers emerging for each task difficulty level. This insight supports the implementation of difficulty-aware evaluators [2024arXiv240407720S, park2024large, dutulescu2024hard, xu2024adaption] that dynamically route queries to the most energy-efficient model achieving optimal accuracy. Optimal energy consumption $E_{o}$ of an oracle evaluator and realistic energy consumption $E_{r}$ of an imperfect evaluator in practical settings employing a high-performing model as a fallback can be calculated as
where $p_{i}$ denotes the probability that the evaluator correctly predicts difficulty level $i$, $q$ represents the probability of incorrect predictions, and $e_{m_{D}}$ is the energy consumed by the largest, most accurate fallback model. Compared against the baseline approach ($E_{max}=Ne_{m_{D}}$), our analysis shows significant potential energy savings. For instance, using an oracle evaluator, LiveCodeBench and MATH500 benchmarks consumed only 1.75 kWh compared to baseline consumptions of 2.03 kWh and 2.20 kWh, respectively.
Length-wise Early Exit
Figure 5 indicates that correct responses typically require fewer output tokens than incorrect responses, suggesting another promising avenue for efficiency improvements. Traditionally, early-exit methods focus on exiting at certain model layers [del2023skipdecode, chen2023ee, teerapittayanon2016branchynet, zhou2020bert]. Extending these techniques to include token-length threshold will enable models to cease generation when answers become evidently flawed or excessively verbose. Such a mechanism effectively reduces unnecessary computations and subsequent energy consumption. Specifically, Figure 5 illustrates that the correct answers use less than 12000 tokens excluding outliers. We can apply this observation on Figure 4, cutting the generation at the fourth point (13108 tokens) on the RT lines. This gives us 14% accuracy saving without loss of accuracy.
Limitations and Future Work
This study focused solely on operational energy demands during model inference. We did not account for embodied energy from hardware production, deployment, or training large models. We also did not include energy measurements from GPU interconnects V2 (due to limitations of available measurement tools [nvml]) or ancillary hardware components (e.g., CPUs, DRAM, SSDs). We excluded incidental operational overhead, such as query scheduling, and did not measure the energy consumption of alternative hardware architectures [2023arXiv230403208D, 10.1145/3470496.3527405, 10.1145/3140659.3080246, 10.1145/3079856.3080246]. Measuring the impact on energy-accuracy trade-off of optimization works such as speculative decoding or quantization is beyond the scope of this work. Future work should address these gaps to build a comprehensive understanding of energy consumption across the complete lifecycle and deployment environment of LLMs.
@inproceedingsmath500, author = Hendrycks, Dan and Burns, Collin and Kadavath, Saurav and Arora, Akul and Basart, Steven and Tang, Eric and Song, Dawn and Steinhardt, Jacob, booktitle = Proceedings of the Neural Information Processing Systems Track on Datasets and Benchmarks, editor = J. Vanschoren and S. Yeung, pages = , title = Measuring Mathematical Problem Solving With the MATH Dataset, url = https://datasets-benchmarks-proceedings.neurips.cc/paper_files/paper/2021/file/be83ab3ecd0db773eb2dc1b0a17836a1-Paper-round2.pdf, volume = 1, year = 2021
@inproceedingsgpqa, title=Gpqa: A graduate-level google-proof q&a benchmark, author=Rein, David and Hou, Betty Li and Stickland, Asa Cooper and Petty, Jackson and Pang, Richard Yuanzhe and Dirani, Julien and Michael, Julian and Bowman, Samuel R, booktitle=First Conference on Language Modeling, year=2024
@ARTICLEhumaneval, author = Chen, Mark and Tworek, Jerry and Jun, Heewoo and Yuan, Qiming and Ponde de Oliveira Pinto, Henrique and Kaplan, Jared and Edwards, Harri and Burda, Yuri and Joseph, Nicholas and Brockman, Greg and Ray, Alex and Puri, Raul and Krueger, Gretchen and Petrov, Michael and Khlaaf, Heidy and Sastry, Girish and Mishkin, Pamela and Chan, Brooke and Gray, Scott and Ryder, Nick and Pavlov, Mikhail and Power, Alethea and Kaiser, Lukasz and Bavarian, Mohammad and Winter, Clemens and Tillet, Philippe and Petroski Such, Felipe and Cummings, Dave and Plappert, Matthias and Chantzis, Fotios and Barnes, Elizabeth and Herbert-Voss, Ariel and Hebgen Guss, William and Nichol, Alex and Paino, Alex and Tezak, Nikolas and Tang, Jie and Babuschkin, Igor and Balaji, Suchir and Jain, Shantanu and Saunders, William and Hesse, Christopher and Carr, Andrew N. and Leike, Jan and Achiam, Josh and Misra, Vedant and Morikawa, Evan and Radford, Alec and Knight, Matthew and Brundage, Miles and Murati, Mira and Mayer, Katie and Welinder, Peter and McGrew, Bob and Amodei, Dario and McCandlish, Sam and Sutskever, Ilya and Zaremba, Wojciech, title = "Evaluating Large Language Models Trained on Code", journal = arXiv e-prints, keywords = Computer Science - Machine Learning, year = 2021, month = jul, eid = arXiv:2107.03374, pages = arXiv:2107.03374, doi = 10.48550/arXiv.2107.03374, archivePrefix = arXiv, eprint = 2107.03374, primaryClass = cs.LG, adsurl = https://ui.adsabs.harvard.edu/abs/2021arXiv210703374C, adsnote = Provided by the SAO/NASA Astrophysics Data System
@ARTICLElivecodebench, author = Jain, Naman and Han, King and Gu, Alex and Li, Wen-Ding and Yan, Fanjia and Zhang, Tianjun and Wang, Sida and Solar-Lezama, Armando and Sen, Koushik and Stoica, Ion, title = "LiveCodeBench: Holistic and Contamination Free Evaluation of Large Language Models for Code", journal = arXiv e-prints, keywords = Computer Science - Software Engineering, Computer Science - Computation and Language, Computer Science - Machine Learning, year = 2024, month = mar, eid = arXiv:2403.07974, pages = arXiv:2403.07974, doi = 10.48550/arXiv.2403.07974, archivePrefix = arXiv, eprint = 2403.07974, primaryClass = cs.SE, adsurl = https://ui.adsabs.harvard.edu/abs/2024arXiv240307974J, adsnote = Provided by the SAO/NASA Astrophysics Data System
@ARTICLEmbpp, author = Austin, Jacob and Odena, Augustus and Nye, Maxwell and Bosma, Maarten and Michalewski, Henryk and Dohan, David and Jiang, Ellen and Cai, Carrie and Terry, Michael and Le, Quoc and Sutton, Charles, title = "Program Synthesis with Large Language Models", journal = arXiv e-prints, keywords = Computer Science - Programming Languages, Computer Science - Machine Learning, year = 2021, month = aug, eid = arXiv:2108.07732, pages = arXiv:2108.07732, doi = 10.48550/arXiv.2108.07732, archivePrefix = arXiv, eprint = 2108.07732, primaryClass = cs.PL, adsurl = https://ui.adsabs.harvard.edu/abs/2021arXiv210807732A, adsnote = Provided by the SAO/NASA Astrophysics Data System
@ARTICLEhellaswag, author = Zellers, Rowan and Holtzman, Ari and Bisk, Yonatan and Farhadi, Ali and Choi, Yejin, title = "HellaSwag: Can a Machine Really Finish Your Sentence?", journal = arXiv e-prints, keywords = Computer Science - Computation and Language, year = 2019, month = may, eid = arXiv:1905.07830, pages = arXiv:1905.07830, doi = 10.48550/arXiv.1905.07830, archivePrefix = arXiv, eprint = 1905.07830, primaryClass = cs.CL, adsurl = https://ui.adsabs.harvard.edu/abs/2019arXiv190507830Z, adsnote = Provided by the SAO/NASA Astrophysics Data System
@ARTICLEgsm8k, author = Cobbe, Karl and Kosaraju, Vineet and Bavarian, Mohammad and Chen, Mark and Jun, Heewoo and Kaiser, Lukasz and Plappert, Matthias and Tworek, Jerry and Hilton, Jacob and Nakano, Reiichiro and Hesse, Christopher and Schulman, John, title = "Training Verifiers to Solve Math Word Problems", journal = arXiv e-prints, keywords = Computer Science - Machine Learning, Computer Science - Computation and Language, year = 2021, month = oct, eid = arXiv:2110.14168, pages = arXiv:2110.14168, doi = 10.48550/arXiv.2110.14168, archivePrefix = arXiv, eprint = 2110.14168, primaryClass = cs.LG, adsurl = https://ui.adsabs.harvard.edu/abs/2021arXiv211014168C, adsnote = Provided by the SAO/NASA Astrophysics Data System
@articlemmlu, title=Measuring Massive Multitask Language Understanding, author=Dan Hendrycks and Collin Burns and Steven Basart and Andy Zou and Mantas Mazeika and Dawn Song and Jacob Steinhardt, journal=Proceedings of the International Conference on Learning Representations (ICLR), year=2021
@misca100, author = NVIDIA, title = NVIDIA A100 GPU, url = https://www.nvidia.com/content/dam/en-zz/Solutions/Data-Center/a100/pdf/nvidia-a100-datasheet-nvidia-us-2188504-web.pdf, year = 2020
@miscnvml, author = NVIDIA, title = NVIDIA NVML Documentation, url = https://developer.nvidia.com/management-library-nvml, year = 2024
@ARTICLE2023arXiv231001798H, author = Huang, Jie and Chen, Xinyun and Mishra, Swaroop and Zheng, Huaixiu Steven and Yu, Adams Wei and Song, Xinying and Zhou, Denny, title = "Large Language Models Cannot Self-Correct Reasoning Yet", journal = arXiv e-prints, keywords = Computer Science - Computation and Language, Computer Science - Artificial Intelligence, year = 2023, month = oct, eid = arXiv:2310.01798, pages = arXiv:2310.01798, doi = 10.48550/arXiv.2310.01798, archivePrefix = arXiv, eprint = 2310.01798, primaryClass = cs.CL, adsurl = https://ui.adsabs.harvard.edu/abs/2023arXiv231001798H, adsnote = Provided by the SAO/NASA Astrophysics Data System
@ARTICLEdeepseek, author = DeepSeek-AI and Guo, Daya and Yang, Dejian and Zhang, Haowei and Song, Junxiao and Zhang, Ruoyu and Xu, Runxin and Zhu, Qihao and Ma, Shirong and Wang, Peiyi and Bi, Xiao and Zhang, Xiaokang and Yu, Xingkai and Wu, Yu and Wu, Z. F. and Gou, Zhibin and Shao, Zhihong and Li, Zhuoshu and Gao, Ziyi and Liu, Aixin and Xue, Bing and Wang, Bingxuan and Wu, Bochao and Feng, Bei and Lu, Chengda and Zhao, Chenggang and Deng, Chengqi and Zhang, Chenyu and Ruan, Chong and Dai, Damai and Chen, Deli and Ji, Dongjie and Li, Erhang and Lin, Fangyun and Dai, Fucong and Luo, Fuli and Hao, Guangbo and Chen, Guanting and Li, Guowei and Zhang, H. and Bao, Han and Xu, Hanwei and Wang, Haocheng and Ding, Honghui and Xin, Huajian and Gao, Huazuo and Qu, Hui and Li, Hui and Guo, Jianzhong and Li, Jiashi and Wang, Jiawei and Chen, Jingchang and Yuan, Jingyang and Qiu, Junjie and Li, Junlong and Cai, J. L. and Ni, Jiaqi and Liang, Jian and Chen, Jin and Dong, Kai and Hu, Kai and Gao, Kaige and Guan, Kang and Huang, Kexin and Yu, Kuai and Wang, Lean and Zhang, Lecong and Zhao, Liang and Wang, Litong and Zhang, Liyue and Xu, Lei and Xia, Leyi and Zhang, Mingchuan and Zhang, Minghua and Tang, Minghui and Li, Meng and Wang, Miaojun and Li, Mingming and Tian, Ning and Huang, Panpan and Zhang, Peng and Wang, Qiancheng and Chen, Qinyu and Du, Qiushi and Ge, Ruiqi and Zhang, Ruisong and Pan, Ruizhe and Wang, Runji and Chen, R. J. and Jin, R. L. and Chen, Ruyi and Lu, Shanghao and Zhou, Shangyan and Chen, Shanhuang and Ye, Shengfeng and Wang, Shiyu and Yu, Shuiping and Zhou, Shunfeng and Pan, Shuting and Li, S. S. and Zhou, Shuang and Wu, Shaoqing and Ye, Shengfeng and Yun, Tao and Pei, Tian and Sun, Tianyu and Wang, T. and Zeng, Wangding and Zhao, Wanjia and Liu, Wen and Liang, Wenfeng and Gao, Wenjun and Yu, Wenqin and Zhang, Wentao and Xiao, W. L. and An, Wei and Liu, Xiaodong and Wang, Xiaohan and Chen, Xiaokang and Nie, Xiaotao and Cheng, Xin and Liu, Xin and Xie, Xin and Liu, Xingchao and Yang, Xinyu and Li, Xinyuan and Su, Xuecheng and Lin, Xuheng and Li, X. Q. and Jin, Xiangyue and Shen, Xiaojin and Chen, Xiaosha and Sun, Xiaowen and Wang, Xiaoxiang and Song, Xinnan and Zhou, Xinyi and Wang, Xianzu and Shan, Xinxia and Li, Y. K. and Wang, Y. Q. and Wei, Y. X. and Zhang, Yang and Xu, Yanhong and Li, Yao and Zhao, Yao and Sun, Yaofeng and Wang, Yaohui and Yu, Yi and Zhang, Yichao and Shi, Yifan and Xiong, Yiliang and He, Ying and Piao, Yishi and Wang, Yisong and Tan, Yixuan and Ma, Yiyang and Liu, Yiyuan and Guo, Yongqiang and Ou, Yuan and Wang, Yuduan and Gong, Yue and Zou, Yuheng and He, Yujia and Xiong, Yunfan and Luo, Yuxiang and You, Yuxiang and Liu, Yuxuan and Zhou, Yuyang and Zhu, Y. X. and Xu, Yanhong and Huang, Yanping and Li, Yaohui and Zheng, Yi and Zhu, Yuchen and Ma, Yunxian and Tang, Ying and Zha, Yukun and Yan, Yuting and Ren, Z. Z. and Ren, Zehui and Sha, Zhangli and Fu, Zhe and Xu, Zhean and Xie, Zhenda and Zhang, Zhengyan and Hao, Zhewen and Ma, Zhicheng and Yan, Zhigang and Wu, Zhiyu and Gu, Zihui and Zhu, Zijia and Liu, Zijun and Li, Zilin and Xie, Ziwei and Song, Ziyang and Pan, Zizheng and Huang, Zhen and Xu, Zhipeng and Zhang, Zhongyu and Zhang, Zhen, title = "DeepSeek-R1: Incentivizing Reasoning Capability in LLMs via Reinforcement Learning", journal = arXiv e-prints, keywords = Computer Science - Computation and Language, Computer Science - Artificial Intelligence, Computer Science - Machine Learning, year = 2025, month = jan, eid = arXiv:2501.12948, pages = arXiv:2501.12948, doi = 10.48550/arXiv.2501.12948, archivePrefix = arXiv, eprint = 2501.12948, primaryClass = cs.CL, adsurl = https://ui.adsabs.harvard.edu/abs/2025arXiv250112948D, adsnote = Provided by the SAO/NASA Astrophysics Data System
@ARTICLEllmonk, author = Brown, Bradley and Juravsky, Jordan and Ehrlich, Ryan and Clark, Ronald and Le, Quoc V. and Ré, Christopher and Mirhoseini, Azalia, title = "Large Language Monkeys: Scaling Inference Compute with Repeated Sampling", journal = arXiv e-prints, keywords = Computer Science - Machine Learning, Computer Science - Artificial Intelligence, year = 2024, month = jul, eid = arXiv:2407.21787, pages = arXiv:2407.21787, doi = 10.48550/arXiv.2407.21787, archivePrefix = arXiv, eprint = 2407.21787, primaryClass = cs.LG, adsurl = https://ui.adsabs.harvard.edu/abs/2024arXiv240721787B, adsnote = Provided by the SAO/NASA Astrophysics Data System
@inproceedings orca, author = Gyeong-In Yu and Joo Seong Jeong and Geon-Woo Kim and Soojeong Kim and Byung-Gon Chun, title = Orca: A Distributed Serving System for Transformer-Based Generative Models, booktitle = 16th USENIX Symposium on Operating Systems Design and Implementation (OSDI 22), year = 2022, isbn = 978-1-939133-28-1, address = Carlsbad, CA, pages = 521–538, url = https://www.usenix.org/conference/osdi22/presentation/yu, publisher = USENIX Association, month = jul
@inproceedingsvllm, title=Efficient Memory Management for Large Language Model Serving with PagedAttention, author=Woosuk Kwon and Zhuohan Li and Siyuan Zhuang and Ying Sheng and Lianmin Zheng and Cody Hao Yu and Joseph E. Gonzalez and Hao Zhang and Ion Stoica, booktitle=Proceedings of the ACM SIGOPS 29th Symposium on Operating Systems Principles, year=2023
@inproceedingssglang, author = Zheng, Lianmin and Yin, Liangsheng and Xie, Zhiqiang and Sun, Chuyue and Huang, Jeff and Yu, Cody Hao and Cao, Shiyi and Kozyrakis, Christos and Stoica, Ion and Gonzalez, Joseph E. and Barrett, Clark and Sheng, Ying, booktitle = Advances in Neural Information Processing Systems, editor = A. Globerson and L. Mackey and D. Belgrave and A. Fan and U. Paquet and J. Tomczak and C. Zhang, pages = 62557–62583, publisher = Curran Associates, Inc., title = SGLang: Efficient Execution of Structured Language Model Programs, url = https://proceedings.neurips.cc/paper_files/paper/2024/file/724be4472168f31ba1c9ac630f15dec8-Paper-Conference.pdf, volume = 37, year = 2024
@articlenanoflow, title=Nanoflow: Towards optimal large language model serving throughput, author=Zhu, Kan and Zhao, Yilong and Zhao, Liangyu and Zuo, Gefei and Gu, Yile and Xie, Dedong and Gao, Yufei and Xu, Qinyu and Tang, Tian and Ye, Zihao and others, journal=arXiv preprint arXiv:2408.12757, year=2024
@miscaime24, title=AIME24, author=MAA, url=https://maa.org/math -competitions/american-invitational-mathematics-examination-aime, journal=American invitational mathematics examination - aime. In American Invitational Mathematics Examination - AIME 2024, year=February 2024
@miscaime25, title=AIME25, author=MAA, url=https://maa.org/math -competitions/american-invitational-mathematics-examination-aime, journal=American invitational mathematics examination - aime. In American Invitational Mathematics Examination - AIME 2025, year=February 2024
@articleaiw, title=Alice in Wonderland: Simple Tasks Showing Complete Reasoning Breakdown in State-Of-the-Art Large Language Models, author=Marianna Nezhurina and Lucia Cipolina-Kun and Mehdi Cherti and Jenia Jitsev, year=2024, journal=arXiv preprint arXiv:2406.02061, eprint=2406.02061, archivePrefix=arXiv, primaryClass=cs.LG
@ARTICLEcodeforces, author = Quan, Shanghaoran and Yang, Jiaxi and Yu, Bowen and Zheng, Bo and Liu, Dayiheng and Yang, An and Ren, Xuancheng and Gao, Bofei and Miao, Yibo and Feng, Yunlong and Wang, Zekun and Yang, Jian and Cui, Zeyu and Fan, Yang and Zhang, Yichang and Hui, Binyuan and Lin, Junyang, title = "CodeElo: Benchmarking Competition-level Code Generation of LLMs with Human-comparable Elo Ratings", journal = arXiv e-prints, keywords = Computer Science - Computation and Language, year = 2025, month = jan, eid = arXiv:2501.01257, pages = arXiv:2501.01257, doi = 10.48550/arXiv.2501.01257, archivePrefix = arXiv, eprint = 2501.01257, primaryClass = cs.CL, adsurl = https://ui.adsabs.harvard.edu/abs/2025arXiv250101257Q, adsnote = Provided by the SAO/NASA Astrophysics Data System
@inproceedingsdeepspeed, title=Deepspeed-inference: enabling efficient inference of transformer models at unprecedented scale, author=Aminabadi, Reza Yazdani and Rajbhandari, Samyam and Awan, Ammar Ahmad and Li, Cheng and Li, Du and Zheng, Elton and Ruwase, Olatunji and Smith, Shaden and Zhang, Minjia and Rasley, Jeff and others, booktitle=SC22: International Conference for High Performance Computing, Networking, Storage and Analysis, pages=1–15, year=2022, organization=IEEE
@INPROCEEDINGS9307857, author=Anzt, Hartwig and Tsai, Yuhsiang M. and Abdelfattah, Ahmad and Cojean, Terry and Dongarra, Jack, booktitle=2020 IEEE/ACM Performance Modeling, Benchmarking and Simulation of High Performance Computer Systems (PMBS), title=Evaluating the Performance of NVIDIA’s A100 Ampere GPU for Sparse and Batched Computations, year=2020, volume=, number=, pages=26-38, keywords=Kernel;Graphics processing units;Sparse matrices;Bandwidth;Linear algebra;Libraries;Benchmark testing;Sparse Linear Algebra;Sparse Matrix Vector Product;Batched Linear Algebra;NVIDIA A100 GPU, doi=10.1109/PMBS51919.2020.00009
@miscmath500_leaderboard, author = vals.ai, title = vals.ai MATH500 Leaderboard, url = https://www.vals.ai/benchmarks/math500-03-24-2025, year = 2025
@ARTICLE2024arXiv240407720S, author = Säuberli, Andreas and Clematide, Simon, title = "Automatic Generation and Evaluation of Reading Comprehension Test Items with Large Language Models", journal = arXiv e-prints, keywords = Computer Science - Computation and Language, year = 2024, month = apr, eid = arXiv:2404.07720, pages = arXiv:2404.07720, doi = 10.48550/arXiv.2404.07720, archivePrefix = arXiv, eprint = 2404.07720, primaryClass = cs.CL, adsurl = https://ui.adsabs.harvard.edu/abs/2024arXiv240407720S, adsnote = Provided by the SAO/NASA Astrophysics Data System
@ARTICLEqwen2.5, author = Qwen and : and Yang, An and Yang, Baosong and Zhang, Beichen and Hui, Binyuan and Zheng, Bo and Yu, Bowen and Li, Chengyuan and Liu, Dayiheng and Huang, Fei and Wei, Haoran and Lin, Huan and Yang, Jian and Tu, Jianhong and Zhang, Jianwei and Yang, Jianxin and Yang, Jiaxi and Zhou, Jingren and Lin, Junyang and Dang, Kai and Lu, Keming and Bao, Keqin and Yang, Kexin and Yu, Le and Li, Mei and Xue, Mingfeng and Zhang, Pei and Zhu, Qin and Men, Rui and Lin, Runji and Li, Tianhao and Tang, Tianyi and Xia, Tingyu and Ren, Xingzhang and Ren, Xuancheng and Fan, Yang and Su, Yang and Zhang, Yichang and Wan, Yu and Liu, Yuqiong and Cui, Zeyu and Zhang, Zhenru and Qiu, Zihan, title = "Qwen2.5 Technical Report", journal = arXiv e-prints, keywords = Computer Science - Computation and Language, year = 2024, month = dec, eid = arXiv:2412.15115, pages = arXiv:2412.15115, doi = 10.48550/arXiv.2412.15115, archivePrefix = arXiv, eprint = 2412.15115, primaryClass = cs.CL, adsurl = https://ui.adsabs.harvard.edu/abs/2024arXiv241215115Q, adsnote = Provided by the SAO/NASA Astrophysics Data System
@miscgetting_50, author = Ryan Greenblatt, title = Geting 50, url = https://www.lesswrong.com/posts/Rdwui3wHxCeKb7feK/getting-50-sota-on-arc-agi-with-gpt-4o, year = 2024
@ARTICLE2023arXiv230306135I, author = Irvine, Robert and Boubert, Douglas and Raina, Vyas and Liusie, Adian and Zhu, Ziyi and Mudupalli, Vineet and Korshuk, Aliaksei and Liu, Zongyi and Cremer, Fritz and Assassi, Valentin and Beauchamp, Christie-Carol and Lu, Xiaoding and Rialan, Thomas and Beauchamp, William, title = "Rewarding Chatbots for Real-World Engagement with Millions of Users", journal = arXiv e-prints, keywords = Computer Science - Computation and Language, Computer Science - Artificial Intelligence, Computer Science - Machine Learning, year = 2023, month = mar, eid = arXiv:2303.06135, pages = arXiv:2303.06135, doi = 10.48550/arXiv.2303.06135, archivePrefix = arXiv, eprint = 2303.06135, primaryClass = cs.CL, adsurl = https://ui.adsabs.harvard.edu/abs/2023arXiv230306135I, adsnote = Provided by the SAO/NASA Astrophysics Data System
@ARTICLE2022arXiv220311171W, author = Wang, Xuezhi and Wei, Jason and Schuurmans, Dale and Le, Quoc and Chi, Ed and Narang, Sharan and Chowdhery, Aakanksha and Zhou, Denny, title = "Self-Consistency Improves Chain of Thought Reasoning in Language Models", journal = arXiv e-prints, keywords = Computer Science - Computation and Language, Computer Science - Artificial Intelligence, year = 2022, month = mar, eid = arXiv:2203.11171, pages = arXiv:2203.11171, doi = 10.48550/arXiv.2203.11171, archivePrefix = arXiv, eprint = 2203.11171, primaryClass = cs.CL, adsurl = https://ui.adsabs.harvard.edu/abs/2022arXiv220311171W, adsnote = Provided by the SAO/NASA Astrophysics Data System
@ARTICLE2022arXiv221104325V, author = Villalobos, Pablo and Ho, Anson and Sevilla, Jaime and Besiroglu, Tamay and Heim, Lennart and Hobbhahn, Marius, title = "Will we run out of data? Limits of LLM scaling based on human-generated data", journal = arXiv e-prints, keywords = Computer Science - Machine Learning, Computer Science - Artificial Intelligence, Computer Science - Computation and Language, Computer Science - Computer Vision and Pattern Recognition, Computer Science - Computers and Society, year = 2022, month = oct, eid = arXiv:2211.04325, pages = arXiv:2211.04325, doi = 10.48550/arXiv.2211.04325, archivePrefix = arXiv, eprint = 2211.04325, primaryClass = cs.LG, adsurl = https://ui.adsabs.harvard.edu/abs/2022arXiv221104325V, adsnote = Provided by the SAO/NASA Astrophysics Data System
@techreportshehabi2024united, title = 2024 United States Data Center Energy Usage Report, author = Shehabi, Arman and Smith, Sarah Josephine and Hubbard, Alex and Newkirk, Alexander and Lei, Nuoa and Siddik, Md AbuBakar and Holecek, Billie and Koomey, Jonathan G and Masanet, Eric R and Sartor, Dale A, institution = Lawrence Berkeley National Laboratory, type = Report, number = LBNL-2001637, year = 2024, month = dec, doi = 10.71468/P1WC7Q, url = https://eta-publications.lbl.gov/sites/default/files/2024-12/lbnl-2024-united-states-data-center-energy-usage-report.pdf,
@ARTICLE2023arXiv230510601Y, author = Yao, Shunyu and Yu, Dian and Zhao, Jeffrey and Shafran, Izhak and Griffiths, Thomas L. and Cao, Yuan and Narasimhan, Karthik, title = "Tree of Thoughts: Deliberate Problem Solving with Large Language Models", journal = arXiv e-prints, keywords = Computer Science - Computation and Language, Computer Science - Artificial Intelligence, Computer Science - Machine Learning, year = 2023, month = may, eid = arXiv:2305.10601, pages = arXiv:2305.10601, doi = 10.48550/arXiv.2305.10601, archivePrefix = arXiv, eprint = 2305.10601, primaryClass = cs.CL, adsurl = https://ui.adsabs.harvard.edu/abs/2023arXiv230510601Y, adsnote = Provided by the SAO/NASA Astrophysics Data System
@ARTICLE2023arXiv230317651M, author = Madaan, Aman and Tandon, Niket and Gupta, Prakhar and Hallinan, Skyler and Gao, Luyu and Wiegreffe, Sarah and Alon, Uri and Dziri, Nouha and Prabhumoye, Shrimai and Yang, Yiming and Gupta, Shashank and Prasad Majumder, Bodhisattwa and Hermann, Katherine and Welleck, Sean and Yazdanbakhsh, Amir and Clark, Peter, title = "Self-Refine: Iterative Refinement with Self-Feedback", journal = arXiv e-prints, keywords = Computer Science - Computation and Language, Computer Science - Artificial Intelligence, Computer Science - Machine Learning, year = 2023, month = mar, eid = arXiv:2303.17651, pages = arXiv:2303.17651, doi = 10.48550/arXiv.2303.17651, archivePrefix = arXiv, eprint = 2303.17651, primaryClass = cs.CL, adsurl = https://ui.adsabs.harvard.edu/abs/2023arXiv230317651M, adsnote = Provided by the SAO/NASA Astrophysics Data System
@ARTICLE2024arXiv240615673L, author = Liu, Dancheng and Nassereldine, Amir and Yang, Ziming and Xu, Chenhui and Hu, Yuting and Li, Jiajie and Kumar, Utkarsh and Lee, Changjae and Qin, Ruiyang and Shi, Yiyu and Xiong, Jinjun, title = "Large Language Models have Intrinsic Self-Correction Ability", journal = arXiv e-prints, keywords = Computer Science - Computation and Language, Computer Science - Artificial Intelligence, year = 2024, month = jun, eid = arXiv:2406.15673, pages = arXiv:2406.15673, doi = 10.48550/arXiv.2406.15673, archivePrefix = arXiv, eprint = 2406.15673, primaryClass = cs.CL, adsurl = https://ui.adsabs.harvard.edu/abs/2024arXiv240615673L, adsnote = Provided by the SAO/NASA Astrophysics Data System
@ARTICLE2024arXiv240601297K, author = Kamoi, Ryo and Zhang, Yusen and Zhang, Nan and Han, Jiawei and Zhang, Rui, title = "When Can LLMs Actually Correct Their Own Mistakes? A Critical Survey of Self-Correction of LLMs", journal = arXiv e-prints, keywords = Computer Science - Computation and Language, year = 2024, month = jun, eid = arXiv:2406.01297, pages = arXiv:2406.01297, doi = 10.48550/arXiv.2406.01297, archivePrefix = arXiv, eprint = 2406.01297, primaryClass = cs.CL, adsurl = https://ui.adsabs.harvard.edu/abs/2024arXiv240601297K, adsnote = Provided by the SAO/NASA Astrophysics Data System
@articlewu2022sustainable, title=Sustainable ai: Environmental implications, challenges and opportunities, author=Wu, Carole-Jean and Raghavendra, Ramya and Gupta, Udit and Acun, Bilge and Ardalani, Newsha and Maeng, Kiwan and Chang, Gloria and Aga, Fiona and Huang, Jinshi and Bai, Charles and others, journal=Proceedings of Machine Learning and Systems, volume=4, pages=795–813, year=2022
@articlepatterson2022carbon, title=The carbon footprint of machine learning training will plateau, then shrink, author=Patterson, David and Gonzalez, Joseph and Hölzle, Urs and Le, Quoc and Liang, Chen and Munguia, Lluis-Miquel and Rothchild, Daniel and So, David R and Texier, Maud and Dean, Jeff, journal=Computer, volume=55, number=7, pages=18–28, year=2022, publisher=IEEE
@articlebarr2019amazon, title=Amazon ec2 update–inf1 instances with AWS inferentia chips for high performance cost-effective inferencing, author=Barr, Jeff, journal=AWS News Blog, year=2019
@ARTICLEllama, author = Grattafiori, Aaron and Dubey, Abhimanyu and Jauhri, Abhinav and Pandey, Abhinav and Kadian, Abhishek and Al-Dahle, Ahmad and Letman, Aiesha and Mathur, Akhil and Schelten, Alan and Vaughan, Alex and Yang, Amy and Fan, Angela and Goyal, Anirudh and Hartshorn, Anthony and Yang, Aobo and Mitra, Archi and Sravankumar, Archie and Korenev, Artem and Hinsvark, Arthur and Rao, Arun and Zhang, Aston and Rodriguez, Aurelien and Gregerson, Austen and Spataru, Ava and Roziere, Baptiste and Biron, Bethany and Tang, Binh and Chern, Bobbie and Caucheteux, Charlotte and Nayak, Chaya and Bi, Chloe and Marra, Chris and McConnell, Chris and Keller, Christian and Touret, Christophe and Wu, Chunyang and Wong, Corinne and Canton Ferrer, Cristian and Nikolaidis, Cyrus and Allonsius, Damien and Song, Daniel and Pintz, Danielle and Livshits, Danny and Wyatt, Danny and Esiobu, David and Choudhary, Dhruv and Mahajan, Dhruv and Garcia-Olano, Diego and Perino, Diego and Hupkes, Dieuwke and Lakomkin, Egor and AlBadawy, Ehab and Lobanova, Elina and Dinan, Emily and Smith, Eric Michael and Radenovic, Filip and Guzmán, Francisco and Zhang, Frank and Synnaeve, Gabriel and Lee, Gabrielle and Anderson, Georgia Lewis and Thattai, Govind and Nail, Graeme and Mialon, Gregoire and Pang, Guan and Cucurell, Guillem and Nguyen, Hailey and Korevaar, Hannah and Xu, Hu and Touvron, Hugo and Zarov, Iliyan and Arrieta Ibarra, Imanol and Kloumann, Isabel and Misra, Ishan and Evtimov, Ivan and Zhang, Jack and Copet, Jade and Lee, Jaewon and Geffert, Jan and Vranes, Jana and Park, Jason and Mahadeokar, Jay and Shah, Jeet and van der Linde, Jelmer and Billock, Jennifer and Hong, Jenny and Lee, Jenya and Fu, Jeremy and Chi, Jianfeng and Huang, Jianyu and Liu, Jiawen and Wang, Jie and Yu, Jiecao and Bitton, Joanna and Spisak, Joe and Park, Jongsoo and Rocca, Joseph and Johnstun, Joshua and Saxe, Joshua and Jia, Junteng and Vasuden Alwala, Kalyan and Prasad, Karthik and Upasani, Kartikeya and Plawiak, Kate and Li, Ke and Heafield, Kenneth and Stone, Kevin and El-Arini, Khalid and Iyer, Krithika and Malik, Kshitiz and Chiu, Kuenley and Bhalla, Kunal and Lakhotia, Kushal and Rantala-Yeary, Lauren and van der Maaten, Laurens and Chen, Lawrence and Tan, Liang and Jenkins, Liz and Martin, Louis and Madaan, Lovish and Malo, Lubo and Blecher, Lukas and Landzaat, Lukas and de Oliveira, Luke and Muzzi, Madeline and Pasupuleti, Mahesh and Singh, Mannat and Paluri, Manohar and Kardas, Marcin and Tsimpoukelli, Maria and Oldham, Mathew and Rita, Mathieu and Pavlova, Maya and Kambadur, Melanie and Lewis, Mike and Si, Min and Singh, Mitesh Kumar and Hassan, Mona and Goyal, Naman and Torabi, Narjes and Bashlykov, Nikolay and Bogoychev, Nikolay and Chatterji, Niladri and Zhang, Ning and Duchenne, Olivier and Çelebi, Onur and Alrassy, Patrick and Zhang, Pengchuan and Li, Pengwei and Vasic, Petar and Weng, Peter and Bhargava, Prajjwal and Dubal, Pratik and Krishnan, Praveen and Singh Koura, Punit and Xu, Puxin and He, Qing and Dong, Qingxiao and Srinivasan, Ragavan and Ganapathy, Raj and Calderer, Ramon and Silveira Cabral, Ricardo and Stojnic, Robert and Raileanu, Roberta and Maheswari, Rohan and Girdhar, Rohit and Patel, Rohit and Sauvestre, Romain and Polidoro, Ronnie and Sumbaly, Roshan and Taylor, Ross and Silva, Ruan and Hou, Rui and Wang, Rui and Hosseini, Saghar and Chennabasappa, Sahana and Singh, Sanjay and Bell, Sean and Kim, Seohyun Sonia and Edunov, Sergey and Nie, Shaoliang and Narang, Sharan and Raparthy, Sharath and Shen, Sheng and Wan, Shengye and Bhosale, Shruti and Zhang, Shun and Vandenhende, Simon and Batra, Soumya and Whitman, Spencer and Sootla, Sten and Collot, Stephane and Gururangan, Suchin and Borodinsky, Sydney and Herman, Tamar and Fowler, Tara and Sheasha, Tarek and Georgiou, Thomas and Scialom, Thomas and Speckbacher, Tobias, title = "The Llama 3 Herd of Models", journal = arXiv e-prints, keywords = Computer Science - Artificial Intelligence, Computer Science - Computation and Language, Computer Science - Computer Vision and Pattern Recognition, year = 2024, month = jul, eid = arXiv:2407.21783, pages = arXiv:2407.21783, doi = 10.48550/arXiv.2407.21783, archivePrefix = arXiv, eprint = 2407.21783, primaryClass = cs.AI, adsurl = https://ui.adsabs.harvard.edu/abs/2024arXiv240721783G, adsnote = Provided by the SAO/NASA Astrophysics Data System
@articlegemini, title=Gemini: a family of highly capable multimodal models, author=Team, Gemini and Anil, Rohan and Borgeaud, Sebastian and Alayrac, Jean-Baptiste and Yu, Jiahui and Soricut, Radu and Schalkwyk, Johan and Dai, Andrew M and Hauth, Anja and Millican, Katie and others, journal=arXiv preprint arXiv:2312.11805, year=2023
@articlegpt4, title=Gpt-4 technical report, author=Achiam, Josh and Adler, Steven and Agarwal, Sandhini and Ahmad, Lama and Akkaya, Ilge and Aleman, Florencia Leoni and Almeida, Diogo and Altenschmidt, Janko and Altman, Sam and Anadkat, Shyamal and others, journal=arXiv preprint arXiv:2303.08774, year=2023
@ARTICLEfew_shot, author = Brown, Tom B. and Mann, Benjamin and Ryder, Nick and Subbiah, Melanie and Kaplan, Jared and Dhariwal, Prafulla and Neelakantan, Arvind and Shyam, Pranav and Sastry, Girish and Askell, Amanda and Agarwal, Sandhini and Herbert-Voss, Ariel and Krueger, Gretchen and Henighan, Tom and Child, Rewon and Ramesh, Aditya and Ziegler, Daniel M. and Wu, Jeffrey and Winter, Clemens and Hesse, Christopher and Chen, Mark and Sigler, Eric and Litwin, Mateusz and Gray, Scott and Chess, Benjamin and Clark, Jack and Berner, Christopher and McCandlish, Sam and Radford, Alec and Sutskever, Ilya and Amodei, Dario, title = "Language Models are Few-Shot Learners", journal = arXiv e-prints, keywords = Computer Science - Computation and Language, year = 2020, month = may, eid = arXiv:2005.14165, pages = arXiv:2005.14165, doi = 10.48550/arXiv.2005.14165, archivePrefix = arXiv, eprint = 2005.14165, primaryClass = cs.CL, adsurl = https://ui.adsabs.harvard.edu/abs/2020arXiv200514165B, adsnote = Provided by the SAO/NASA Astrophysics Data System
@articlerest, title=Beyond human data: Scaling self-training for problem-solving with language models, author=Singh, Avi and Co-Reyes, John D and Agarwal, Rishabh and Anand, Ankesh and Patil, Piyush and Garcia, Xavier and Liu, Peter J and Harrison, James and Lee, Jaehoon and Xu, Kelvin and others, journal=arXiv preprint arXiv:2312.06585, year=2023
@articlestar, title=Star: Bootstrapping reasoning with reasoning, 2022, author=Zelikman, Eric and Wu, Yuhuai and Mu, Jesse and Goodman, Noah D, journal=URL https://arxiv. org/abs/2203.14465, volume=2203, year=2022
@inproceedingsstrubell2020energy, title=Energy and policy considerations for modern deep learning research, author=Strubell, Emma and Ganesh, Ananya and McCallum, Andrew, booktitle=Proceedings of the AAAI conference on artificial intelligence, volume=34, pages=13693–13696, year=2020
@articlefernandez2025energy, title=Energy Considerations of Large Language Model Inference and Efficiency Optimizations, author=Fernandez, Jared and Na, Clara and Tiwari, Vashisth and Bisk, Yonatan and Luccioni, Sasha and Strubell, Emma, journal=arXiv preprint arXiv:2504.17674, year=2025
@inproceedingsluccioni2024power, title=Power hungry processing: Watts driving the cost of ai deployment?, author=Luccioni, Sasha and Jernite, Yacine and Strubell, Emma, booktitle=Proceedings of the 2024 ACM conference on fairness, accountability, and transparency, pages=85–99, year=2024
@articlewu2025unveiling, title=Unveiling environmental impacts of large language model serving: A functional unit view, author=Wu, Yanran and Hua, Inez and Ding, Yi, journal=arXiv preprint arXiv:2502.11256, year=2025
@inproceedingspatel2024characterizing, title=Characterizing power management opportunities for llms in the cloud, author=Patel, Pratyush and Choukse, Esha and Zhang, Chaojie and Goiri, Íñigo and Warrier, Brijesh and Mahalingam, Nithish and Bianchini, Ricardo, booktitle=Proceedings of the 29th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 3, pages=207–222, year=2024
@articlestojkovic2024towards, title=Towards greener llms: Bringing energy-efficiency to the forefront of llm inference, author=Stojkovic, Jovan and Choukse, Esha and Zhang, Chaojie and Goiri, Inigo and Torrellas, Josep, journal=arXiv preprint arXiv:2403.20306, year=2024
@articledel2023skipdecode, title=Skipdecode: Autoregressive skip decoding with batching and caching for efficient llm inference, author=Del Corro, Luciano and Del Giorno, Allie and Agarwal, Sahaj and Yu, Bin and Awadallah, Ahmed and Mukherjee, Subhabrata, journal=arXiv preprint arXiv:2307.02628, year=2023
@articlechen2023ee, title=Ee-llm: Large-scale training and inference of early-exit large language models with 3d parallelism, author=Chen, Yanxi and Pan, Xuchen and Li, Yaliang and Ding, Bolin and Zhou, Jingren, journal=arXiv preprint arXiv:2312.04916, year=2023
@inproceedingsteerapittayanon2016branchynet, title=Branchynet: Fast inference via early exiting from deep neural networks, author=Teerapittayanon, Surat and McDanel, Bradley and Kung, Hsiang-Tsung, booktitle=2016 23rd international conference on pattern recognition (ICPR), pages=2464–2469, year=2016, organization=IEEE
@articlezhou2020bert, title=Bert loses patience: Fast and robust inference with early exit, author=Zhou, Wangchunshu and Xu, Canwen and Ge, Tao and McAuley, Julian and Xu, Ke and Wei, Furu, journal=Advances in Neural Information Processing Systems, volume=33, pages=18330–18341, year=2020
@inproceedingspark2024large, title=Large Language Models are Students at Various Levels: Zero-shot Question Difficulty Estimation, author=Park, Jae-Woo and Park, Seong-Jin and Won, Hyun-Sik and Kim, Kang-Min, booktitle=Findings of the Association for Computational Linguistics: EMNLP 2024, pages=8157–8177, year=2024
@inproceedingsdutulescu2024hard, title=How Hard Can This Question Be? An Exploratory Analysis of Features Assessing Question Difficulty Using LLMs, author=Dutulescu, Andreea and Ruseti, Stefan and Dascalu, Mihai and Mcnamara, Danielle, booktitle=Proceedings of the 17th International Conference on Educational Data Mining, pages=802–808, year=2024
@inproceedingsxu2024adaption, title=Adaption-of-thought: Learning question difficulty improves large language models for reasoning, author=Xu, Mayi and Li, Yongqi and Sun, Ke and Qian, Tieyun, booktitle=Proceedings of the 2024 Conference on Empirical Methods in Natural Language Processing, pages=5468–5495, year=2024
@inproceedingstalmor-etal-2019-commonsenseqa, title = "CommonsenseQA: A Question Answering Challenge Targeting Commonsense Knowledge", author = "Talmor, Alon and Herzig, Jonathan and Lourie, Nicholas and Berant, Jonathan", editor = "Burstein, Jill and Doran, Christy and Solorio, Thamar", booktitle = "Proceedings of the 2019 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 1 (Long and Short Papers)", month = jun, year = "2019", address = "Minneapolis, Minnesota", publisher = "Association for Computational Linguistics", url = "https://aclanthology.org/N19-1421/", doi = "10.18653/v1/N19-1421", pages = "4149–4158", abstract = "When answering a question, people often draw upon their rich world knowledge in addition to the particular context. Recent work has focused primarily on answering questions given some relevant document or context, and required very little general background. To investigate question answering with prior knowledge, we present CommonsenseQA: a challenging new dataset for commonsense question answering. To capture common sense beyond associations, we extract from ConceptNet (Speer et al., 2017) multiple target concepts that have the same semantic relation to a single source concept. Crowd-workers are asked to author multiple-choice questions that mention the source concept and discriminate in turn between each of the target concepts. This encourages workers to create questions with complex semantics that often require prior knowledge. We create 12,247 questions through this procedure and demonstrate the difficulty of our task with a large number of strong baselines. Our best baseline is based on BERT-large (Devlin et al., 2018) and obtains 56% accuracy, well below human performance, which is 89%."
@ARTICLE2020arXiv200108361K, author = Kaplan, Jared and McCandlish, Sam and Henighan, Tom and Brown, Tom B. and Chess, Benjamin and Child, Rewon and Gray, Scott and Radford, Alec and Wu, Jeffrey and Amodei, Dario, title = "Scaling Laws for Neural Language Models", journal = arXiv e-prints, keywords = Computer Science - Machine Learning, Statistics - Machine Learning, year = 2020, month = jan, eid = arXiv:2001.08361, pages = arXiv:2001.08361, doi = 10.48550/arXiv.2001.08361, archivePrefix = arXiv, eprint = 2001.08361, primaryClass = cs.LG, adsurl = https://ui.adsabs.harvard.edu/abs/2020arXiv200108361K, adsnote = Provided by the SAO/NASA Astrophysics Data System
@inproceedings10.1145/3079856.3080246, author = Jouppi, Norman P. and Young, Cliff and Patil, Nishant and Patterson, David and Agrawal, Gaurav and Bajwa, Raminder and Bates, Sarah and Bhatia, Suresh and Boden, Nan and Borchers, Al and Boyle, Rick and Cantin, Pierre-luc and Chao, Clifford and Clark, Chris and Coriell, Jeremy and Daley, Mike and Dau, Matt and Dean, Jeffrey and Gelb, Ben and Ghaemmaghami, Tara Vazir and Gottipati, Rajendra and Gulland, William and Hagmann, Robert and Ho, C. Richard and Hogberg, Doug and Hu, John and Hundt, Robert and Hurt, Dan and Ibarz, Julian and Jaffey, Aaron and Jaworski, Alek and Kaplan, Alexander and Khaitan, Harshit and Killebrew, Daniel and Koch, Andy and Kumar, Naveen and Lacy, Steve and Laudon, James and Law, James and Le, Diemthu and Leary, Chris and Liu, Zhuyuan and Lucke, Kyle and Lundin, Alan and MacKean, Gordon and Maggiore, Adriana and Mahony, Maire and Miller, Kieran and Nagarajan, Rahul and Narayanaswami, Ravi and Ni, Ray and Nix, Kathy and Norrie, Thomas and Omernick, Mark and Penukonda, Narayana and Phelps, Andy and Ross, Jonathan and Ross, Matt and Salek, Amir and Samadiani, Emad and Severn, Chris and Sizikov, Gregory and Snelham, Matthew and Souter, Jed and Steinberg, Dan and Swing, Andy and Tan, Mercedes and Thorson, Gregory and Tian, Bo and Toma, Horia and Tuttle, Erick and Vasudevan, Vijay and Walter, Richard and Wang, Walter and Wilcox, Eric and Yoon, Doe Hyun, title = In-Datacenter Performance Analysis of a Tensor Processing Unit, year = 2017, isbn = 9781450348928, publisher = Association for Computing Machinery, address = New York, NY, USA, url = https://doi.org/10.1145/3079856.3080246, doi = 10.1145/3079856.3080246, abstract = Many architects believe that major improvements in cost-energy-performance must now come from domain-specific hardware. This paper evaluates a custom ASIC—called a Tensor Processing Unit (TPU) — deployed in datacenters since 2015 that accelerates the inference phase of neural networks (NN). The heart of the TPU is a 65,536 8-bit MAC matrix multiply unit that offers a peak throughput of 92 TeraOps/second (TOPS) and a large (28 MiB) software-managed on-chip memory. The TPU’s deterministic execution model is a better match to the 99th-percentile response-time requirement of our NN applications than are the time-varying optimizations of CPUs and GPUs that help average throughput more than guaranteed latency. The lack of such features helps explain why, despite having myriad MACs and a big memory, the TPU is relatively small and low power. We compare the TPU to a server-class Intel Haswell CPU and an Nvidia K80 GPU, which are contemporaries deployed in the same datacenters. Our workload, written in the high-level TensorFlow framework, uses production NN applications (MLPs, CNNs, and LSTMs) that represent 95% of our datacenters’ NN inference demand. Despite low utilization for some applications, the TPU is on average about 15X – 30X faster than its contemporary GPU or CPU, with TOPS/Watt about 30X – 80X higher. Moreover, using the CPU’s GDDR5 memory in the TPU would triple achieved TOPS and raise TOPS/Watt to nearly 70X the GPU and 200X the CPU., booktitle = Proceedings of the 44th Annual International Symposium on Computer Architecture, pages = 1–12, numpages = 12, keywords = CNN, DNN, GPU, LSTM, MLP, RNN, TPU, TensorFlow, accelerator, deep learning, domain-specific architecture, neural network, location = Toronto, ON, Canada, series = ISCA ’17
@article10.1145/3140659.3080246, author = Jouppi, Norman P. and Young, Cliff and Patil, Nishant and Patterson, David and Agrawal, Gaurav and Bajwa, Raminder and Bates, Sarah and Bhatia, Suresh and Boden, Nan and Borchers, Al and Boyle, Rick and Cantin, Pierre-luc and Chao, Clifford and Clark, Chris and Coriell, Jeremy and Daley, Mike and Dau, Matt and Dean, Jeffrey and Gelb, Ben and Ghaemmaghami, Tara Vazir and Gottipati, Rajendra and Gulland, William and Hagmann, Robert and Ho, C. Richard and Hogberg, Doug and Hu, John and Hundt, Robert and Hurt, Dan and Ibarz, Julian and Jaffey, Aaron and Jaworski, Alek and Kaplan, Alexander and Khaitan, Harshit and Killebrew, Daniel and Koch, Andy and Kumar, Naveen and Lacy, Steve and Laudon, James and Law, James and Le, Diemthu and Leary, Chris and Liu, Zhuyuan and Lucke, Kyle and Lundin, Alan and MacKean, Gordon and Maggiore, Adriana and Mahony, Maire and Miller, Kieran and Nagarajan, Rahul and Narayanaswami, Ravi and Ni, Ray and Nix, Kathy and Norrie, Thomas and Omernick, Mark and Penukonda, Narayana and Phelps, Andy and Ross, Jonathan and Ross, Matt and Salek, Amir and Samadiani, Emad and Severn, Chris and Sizikov, Gregory and Snelham, Matthew and Souter, Jed and Steinberg, Dan and Swing, Andy and Tan, Mercedes and Thorson, Gregory and Tian, Bo and Toma, Horia and Tuttle, Erick and Vasudevan, Vijay and Walter, Richard and Wang, Walter and Wilcox, Eric and Yoon, Doe Hyun, title = In-Datacenter Performance Analysis of a Tensor Processing Unit, year = 2017, issue_date = May 2017, publisher = Association for Computing Machinery, address = New York, NY, USA, volume = 45, number = 2, issn = 0163-5964, url = https://doi.org/10.1145/3140659.3080246, doi = 10.1145/3140659.3080246, abstract = Many architects believe that major improvements in cost-energy-performance must now come from domain-specific hardware. This paper evaluates a custom ASIC—called a Tensor Processing Unit (TPU) — deployed in datacenters since 2015 that accelerates the inference phase of neural networks (NN). The heart of the TPU is a 65,536 8-bit MAC matrix multiply unit that offers a peak throughput of 92 TeraOps/second (TOPS) and a large (28 MiB) software-managed on-chip memory. The TPU’s deterministic execution model is a better match to the 99th-percentile response-time requirement of our NN applications than are the time-varying optimizations of CPUs and GPUs that help average throughput more than guaranteed latency. The lack of such features helps explain why, despite having myriad MACs and a big memory, the TPU is relatively small and low power. We compare the TPU to a server-class Intel Haswell CPU and an Nvidia K80 GPU, which are contemporaries deployed in the same datacenters. Our workload, written in the high-level TensorFlow framework, uses production NN applications (MLPs, CNNs, and LSTMs) that represent 95% of our datacenters’ NN inference demand. Despite low utilization for some applications, the TPU is on average about 15X – 30X faster than its contemporary GPU or CPU, with TOPS/Watt about 30X – 80X higher. Moreover, using the CPU’s GDDR5 memory in the TPU would triple achieved TOPS and raise TOPS/Watt to nearly 70X the GPU and 200X the CPU., journal = SIGARCH Comput. Archit. News, month = jun, pages = 1–12, numpages = 12, keywords = CNN, DNN, GPU, LSTM, MLP, RNN, TPU, TensorFlow, accelerator, deep learning, domain-specific architecture, neural network
@inproceedings10.1145/3470496.3527405, author = Abts, Dennis and Kimmell, Garrin and Ling, Andrew and Kim, John and Boyd, Matt and Bitar, Andrew and Parmar, Sahil and Ahmed, Ibrahim and DiCecco, Roberto and Han, David and Thompson, John and Bye, Michael and Hwang, Jennifer and Fowers, Jeremy and Lillian, Peter and Murthy, Ashwin and Mehtabuddin, Elyas and Tekur, Chetan and Sohmers, Thomas and Kang, Kris and Maresh, Stephen and Ross, Jonathan, title = A software-defined tensor streaming multiprocessor for large-scale machine learning, year = 2022, isbn = 9781450386104, publisher = Association for Computing Machinery, address = New York, NY, USA, url = https://doi.org/10.1145/3470496.3527405, doi = 10.1145/3470496.3527405, abstract = We describe our novel commercial software-defined approach for large-scale interconnection networks of tensor streaming processing (TSP) elements. The system architecture includes packaging, routing, and flow control of the interconnection network of TSPs. We describe the communication and synchronization primitives of a bandwidth-rich substrate for global communication. This scalable communication fabric provides the backbone for large-scale systems based on a software-defined Dragonfly topology, ultimately yielding a parallel machine learning system with elasticity to support a variety of workloads, both training and inference. We extend the TSP’s producer-consumer stream programming model to include global memory which is implemented as logically shared, but physically distributed SRAM on-chip memory. Each TSP contributes 220 MiBytes to the global memory capacity, with the maximum capacity limited only by the network’s scale — the maximum number of endpoints in the system. The TSP acts as both a processing element (endpoint) and network switch for moving tensors across the communication links. We describe a novel software-controlled networking approach that avoids the latency variation introduced by dynamic contention for network links. We describe the topology, routing and flow control to characterize the performance of the network that serves as the fabric for a large-scale parallel machine learning system with up to 10,440 TSPs and more than 2 TeraBytes of global memory accessible in less than 3 microseconds of end-to-end system latency., booktitle = Proceedings of the 49th Annual International Symposium on Computer Architecture, pages = 567–580, numpages = 14, keywords = tensor streaming processor, software scheduling, machine learning, dragonfly, location = New York, New York, series = ISCA ’22
@INPROCEEDINGS9731612, author=Prabhakar, Raghu and Jairath, Sumti and Shin, Jinuk Luke, booktitle=2022 IEEE International Solid-State Circuits Conference (ISSCC), title=SambaNova SN10 RDU: A 7nm Dataflow Architecture to Accelerate Software 2.0, year=2022, volume=65, number=, pages=350-352, keywords=Computational modeling;Pipelines;Memory management;Software algorithms;Parallel processing;Software;Natural language processing, doi=10.1109/ISSCC42614.2022.9731612
@ARTICLE2023arXiv230403208D, author = Dey, Nolan and Gosal, Gurpreet and Zhiming and Chen and Khachane, Hemant and Marshall, William and Pathria, Ribhu and Tom, Marvin and Hestness, Joel, title = "Cerebras-GPT: Open Compute-Optimal Language Models Trained on the Cerebras Wafer-Scale Cluster", journal = arXiv e-prints, keywords = Computer Science - Machine Learning, Computer Science - Computation and Language, year = 2023, month = apr, eid = arXiv:2304.03208, pages = arXiv:2304.03208, doi = 10.48550/arXiv.2304.03208, archivePrefix = arXiv, eprint = 2304.03208, primaryClass = cs.LG, adsurl = https://ui.adsabs.harvard.edu/abs/2023arXiv230403208D, adsnote = Provided by the SAO/NASA Astrophysics Data System
@miscevalchemy, author = Guha, Etash and Raoof, Negin and Mercat, Jean and Marten, Ryan and Frankel, Eric and Keh, Sedrick and Grover, Sachin and Smyrnis, George and Vu, Trung and Saad-Falcon, Jon and Choi, Caroline and Arora, Kushal and Merrill, Mike and Deng, Yichuan and Suvarna, Ashima and Bansal, Hritik and Nezhurina, Marianna and Choi, Yejin and Heckel, Reinhard and Oh, Seewong and Hashimoto, Tatsunori and Jitsev, Jenia and Shankar, Vaishaal and Dimakis, Alex and Sathiamoorthy, Mahesh and Schmidt, Ludwig, month = nov, title = Evalchemy: Automatic evals for LLMs, year = 2024